OpenStudy (anonymous):

The function f(x) is shown below. f(x) = (3)^x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? ANSWER CHOICES f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13

OpenStudy (anonymous):

ANSWER CHOICES f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13

OpenStudy (anonymous):

@suckerofmath

OpenStudy (anonymous):

@Justin-the-Nerd

OpenStudy (anonymous):

I'd say C but I'm not sure.

OpenStudy (anonymous):

@suckerofmath oh i forgot, anything after the parenthesis is exponent, so for C it would be f(x) = (3) to the power of x-5

OpenStudy (campbell_st):

well you'll need to look at the indexes what value of x makes x - 2 = 0 Solution and to move it 13 to the left... its 13 - solution = answer then it will be 3^{x + answer} hope it makes sense

OpenStudy (anonymous):

@campbell_st so... C ?

OpenStudy (campbell_st):

nope... the solution to x -2 = 0 is x = 2 now what is 13 - 2 = Answer and the solution is 3^(x + answer)

OpenStudy (anonymous):

3^(x+11) ??

OpenStudy (campbell_st):

thats what my answer would be

OpenStudy (anonymous):

@campbell_st alright, ill choose A then wish me luck, can you help me with another one ?

OpenStudy (campbell_st):

sorry... I have to go and walk the dogs...

OpenStudy (anonymous):

haha alright, have fun